OpenStudy (anonymous):

A square block of steel with volume 10 cm3 and mass of 75 g is cut precisely in half. The density of the two smaller pieces is now...

OpenStudy (anonymous):

a two times the original density. b one-fourth the original density. c the same as the original density.

OpenStudy (anonymous):

a or b not really sure

OpenStudy (anonymous):

\[\begin{align*} \rho &= \frac{m}{V} \\ \\ &= \frac{\dfrac{1}{2}m}{\dfrac{1}{2}V}\end{align*}\] Cutting the block in half reduces both the mass and volume of each piece to half of what they were before. The density of the material does not change.

OpenStudy (anonymous):

so then its A? @ghuczek

OpenStudy (anonymous):

NOT A I MEAN B @ghuczek

OpenStudy (anonymous):

No. As I said in my post, the density does not change. That means it has the be the same as it was before.